# Generating Clinically Realistic EHR Data via a Hierarchy- and Semantics-Guided Transformer

### v1

@@ Abstract
Generating realistic synthetic electronic health records (EHRs) holds tremendous promise for accelerating healthcare research, facilitating AI model development and enhancing patient privacy. However, existing generative methods typically treat EHRs as flat sequences of discrete medical codes. This approach overlooks two critical aspects: the inherent hierarchical organization of clinical coding systems and the rich semantic context provided by code descriptions. Consequently, synthetic patient sequences often lack high clinical fidelity and have limited utility in downstream clinical tasks. In this paper, we propose the Hierarchy- and Semantics-Guided Transformer (HiSGT), a novel framework that leverages both hierarchical and semantic information for the generative process. HiSGT constructs a hierarchical graph to encode parent-child and sibling relationships among clinical codes and employs a graph neural network to derive hierarchy-aware embeddings. These are then fused with semantic embeddings extracted from a pre-trained clinical language model (e.g., ClinicalBERT), enabling the Transformer-based generator to more accurately model the nuanced clinical patterns inherent in real EHRs. Extensive experiments on the MIMIC-III and MIMIC-IV datasets demonstrate that HiSGT significantly improves the statistical alignment of synthetic data with real patient records, as well as supports robust downstream applications such as chronic disease classification. Our evaluations, based on fidelity, utility, and privacy metrics, indicate that HiSGT not only enhances the quality of synthetic EHR data but also preserves patient privacy. By addressing the limitations of conventional raw code-based generative models, HiSGT represents a significant step toward clinically high-fidelity synthetic data generation and a general paradigm suitable for interpretable medical code representation, offering valuable applications in data augmentation and privacy-preserving healthcare analytics.
Keywords Electronic health records  $\cdot$ Synthetic data generation  $\cdot$ Hierarchy and semantics-guided  $\cdot$ Generative transformer
@@ 1 Introduction
Electronic Health Records (EHRs) are a cornerstone of modern healthcare, playing a pivotal role in patient care, clinical decision-making, and medical research (Jha et al., 2009; Cowie et al., 2017; Xiao and Sun, 2021; Xiao et al., 2018). Despite their importance, strict regulations surrounding privacy and data confidentiality have made it challenging to share and utilize EHR data at scale (Johnson et al., 2016, 2023). As a result, there is growing interest in synthetic EHR data generation, a paradigm that aims to produce clinically realistic datasets mimicking the distributional properties of real EHRs while protecting sensitive patient information (Gonzales et al., 2023; McDuff et al., 2023; van Breugel et al., 2024; Ghosheh et al., 2024). By balancing privacy preservation and data utility, synthetic EHRs hold the promise of accelerating model benchmarking, enabling data augmentation, and facilitating collaborative research.
Recent advances in generative models have spurred significant progress in synthetic EHR data generation. Many approaches use Generative Adversarial Networks (GANs) (Choi et al., 2017; Zhang et al., 2021; Torfi and Fox, 2020; Cui et al., 2020; Zhang et al., 2020; Rashidian et al., 2020; Kuo et al., 2022) and Variational AutoEncoders (VAEs) (Biswal et al., 2021). Other methods have drawn on the Transformer architecture (Vaswani et al., 2017) to capture long-range dependencies in clinical sequences. In these models, medical codes and clinical events are typically treated as discrete tokens, with a next-token prediction objective used to model the sequential structure of EHR data (Theodorou et al., 2023; Kraljevic et al., 2024; Pang et al., 2024; Renc et al., 2024). However, they overlook the rich semantic information embedded in clinical code descriptions and the inherent hierarchical relationships in clinical coding systems. Consequently, the synthesized patient sequences frequently lack high clinical fidelity, which limits the utility of the generated synthetic EHR data. In particular, two major research gaps remain.
First, medical coding systems such as the International Classification of Diseases (ICD) (Organization et al., 1978) are organized hierarchically, with codes exhibiting parent–child and sibling relationships that reflect key clinical knowledge. For example, in ICD-10, the code ‘E11’ (Type 2 diabetes mellitus) serves as a parent for ‘E11.6’ (type 2 diabetes mellitus with other specified complications) and as an ancestor for ‘E11.65’ (Type 2 diabetes mellitus with hyperglycemia), while sibling codes such as ‘E11.65’ and ‘E11.64’ (Type 2 diabetes mellitus with hypoglycemia) capture different diabetes complications at a more granular level. In clinical practice, a patient may first be diagnosed at a coarse level with ‘E11’ and later be assessed for specific complications (e.g., ‘E11.64’ or ‘E11.65’). However, current generative models often treat each code as an isolated symbol, failing to leverage these multi-level taxonomic properties. As a result, they may generate conflicting sibling codes that cannot clinically co-occur (e.g., ‘E11.64’ hypoglycemia and ‘E11.65’ hyperglycemia in the same visit) or neglect important subcodes. Second, each medical code is accompanied by a textual description that clarifies its clinical semantics and distinguishes between disease subtypes. For instance, ‘E11.65’ denotes ‘Type 2 diabetes mellitus with hyperglycemia’, while ‘Z13.6’ specifies ‘screening for cardiovascular disorders’. Such text descriptions are essential for determining the precise clinical focus of a patient’s visit. In practice, each patient’s visit to the hospital often centers on a primary diagnosis (or a small set of diagnoses) with additional procedures tailored to that theme until the patient’s condition is confirmed by the physician. Consequently, the codes assigned within a single visit tend to be semantically related, reflecting the shared focus on a particular disease process. However, existing generative methods typically disregard these semantic descriptions and treat codes solely as discrete tokens. Without leveraging the underlying semantics, these models may synthesize codes that are logically or clinically incompatible.
Beyond these challenges, EHR data are inherently high-dimensional and sparse. For example, datasets like MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023) contain thousands of unique diagnostic codes, many of which occur infrequently. This sparsity and high dimensionality pose challenges in capturing the statistical and clinical relationships among codes. Without incorporating additional information—such as textual descriptions or hierarchical structures—purely raw code-based methods struggle with these nuances. As a result, many generative approaches mitigate this issue by reducing data complexity, often through aggregating medical codes or removing rare codes (Theodorou et al., 2023). Consequently, the generated synthetic patient sequences often exhibit low clinical fidelity, ultimately limiting their practical utility in real-world healthcare applications.
@@
To address these limitations, we propose a novel framework called Hierarchy and Semantics-Guided Transformer (HiSGT), which integrates both hierarchical relationships and semantic information into the generative process. HiGST begins by constructing a hierarchical graph to mirror the taxonomy of real coding systems such as ICD, preserving the parent–child and sibling relationships. This graph helps to encode and extract clinically coherent paths for every single code (e.g., from ‘E11.6’ to ‘E11.65’). A graph neural network (GNN) is then employed to learn hierarchy-aware embeddings that capture structured dependencies among medical codes. Additionally, HiGST enriches each medical code with semantic embeddings derived from a pre-trained clinical language model (e.g. ClinicalBERT (Alsentzer et al., 2019)). These embeddings are consistent with their text description to help inject essential domain knowledge into the Transformer backbone. Thus, concepts like ‘Type 2 diabetes’ carry meaningful distinctions rather than being treated solely as a meaningless token. Finally, these refined embeddings are fed into a Transformer-based generative model to synthesize EHR data that are both semantically consistent and structurally coherent. Our work makes the following key contributions to the field of synthetic EHR generation: (1) we propose a novel framework that leverages both hierarchical graph representations and clinical semantic embeddings, as an approach to embed medical codes in general, to overcome the limitations of raw code-only generative methods; (2) we develop an enhanced Transformer model that integrates refined code representations, enabling the generation of synthetic EHRs with high clinical fidelity; (3) we demonstrate the effectiveness of HiSGT on two real-world EHR datasets, showing significant improvements in data quality, downstream task performance (e.g., disease classification), and privacy preservation.
2 Results
2.1 Fidelity Evaluation
To assess the clinical fidelity of synthetic EHR data, we employ four complementary metrics that assess the alignment between real and synthetic records at different levels of granularity, capturing both individual event distributions and their relationships within and across patient visits. First, we compute the Unigram score, which measures how well the generated data preserves the marginal distribution of individual medical events. For each event, we count its occurrences in both the real and synthetic datasets and compute the coefficient $R^{2}$ between the two frequency distributions. A higher $R^{2}$ value indicates that the generated data preserves the marginal distribution of real-world medical events. Beyond individual event distributions, we assess co-occurrence structures by computing the Bigram score, which captures the statistical dependencies between pairs of medical events within the same visit. This score reflects intra-visit coherence—an essential property of high-fidelity EHR synthesis, as medical events within a visit often exhibit strong correlations. Similarly, the sequential bigram score (Seq Bigram) extends this evaluation to inter-visit dependencies, quantifying how well the synthetic data captures longitudinal patterns by examining the transitions between medical codes across consecutive visits. Finally, we compute the dimension-wise correlation (DimWise). Unlike Unigram that only captures overall dataset-level frequency distributions, DimWise measures the per-patient probability distribution of medical codes between real and synthetic data. We first construct code probabilities matrices for real and synthetic datasets, where each row represents a patient and each column represents an ICD code. The elements of the matrix correspond to the relative frequency of each ICD code within a given patient’s sequence. Specifically, for a patient with $n$ total medical codes, the probability of each code is computed by dividing its occurrence count by $n$. We then compute the mean probability of each ICD code across all patients, reducing the original patient-code matrix to a single probability vector per dataset. The coefficient of determination ($R^{2}$ score) is computed between these two vectors (one for real and one for synthetic data) as the final DimWise score.
@@ 2.1 Fidelity Evaluation
To further evaluate the fidelity of synthetic EHR data, we visualize the alignment between real and synthetic medical code distributions using a scatter plot in Figure 2. Each point represents a unique medical code or medical code pair, and its occurrence frequency in the synthetic dataset plotted against its corresponding frequency in real data. Ideally, all points should lie along the diagonal red line, indicating a perfect match between synthetic and real distributions. From the figure, we observe that HiSGT exhibits a significantly tighter clustering of points along the diagonal compared to baseline models, while baseline models such as LSTM, GPT, and SynTEG show greater dispersion. This visualization figure supports the results in Table 1, where HiSGT consistently outperforms competing methods on Unigram, Bigram, and Seq Bigram scores.
Figure 2: Scatter plot of synthetic vs. real distributions for Unigram, Bigram and Seq Bigram on MIMIC-IV. Each point represents a unique medical code or medical code pair, where its frequency in the synthetic dataset is plotted against its frequency in real data. The red diagonal line indicates perfect alignment with real distributions. HiSGT exhibits a tighter clustering along the diagonal compared to baselines, demonstrating superior fidelity in preserving real-world code distributions.
2.2 Utility Evaluation
Table 2: Performance on Phenotype Classification. This table presents the results of the Train on Synthetic, Test on Real (TSTR) evaluation for phenotype classification. Models are trained on synthetic patient records and evaluated on real MIMIC-III and MIMIC-IV datasets. We report the average accuracy, precision, recall, and F1-score (with standard deviations) across 25 phenotype categories. Higher scores indicate better generalization to real-world patient data.
@@ 2.2 Utility Evaluation
Table 3: Phenotype Classification Performance Across Disease Categories. This table provides a detailed breakdown of phenotype classification performance for different disease types (acute, chronic, and mixed). We compare models trained on synthetic data (HALO and HiSGT) with real data as the ground truth. Accuracy and F1-scores are reported for each phenotype category, with macro-averaged scores included for all acute, mixed, and chronic disease groups. Results demonstrate the ability of synthetic data to support downstream predictive modeling across a diverse range of clinical conditions.

### v2

@@ Abstract
Generating realistic synthetic electronic health records (EHRs) holds tremendous promise for accelerating healthcare research, facilitating AI model development and enhancing patient privacy. However, existing generative methods typically treat EHRs as flat sequences of discrete medical codes. This approach overlooks two critical aspects: the inherent hierarchical organization of clinical coding systems and the rich semantic context provided by code descriptions. Consequently, synthetic patient sequences often lack high clinical fidelity and have limited utility in downstream clinical tasks. In this paper, we propose the Hierarchy- and Semantics-Guided Transformer (HiSGT), a novel framework that leverages both hierarchical and semantic information for the generative process. HiSGT constructs a hierarchical graph to encode parent-child and sibling relationships among clinical codes and employs a graph neural network to derive hierarchy-aware embeddings. These are then fused with semantic embeddings extracted from a pre-trained clinical language model (e.g., ClinicalBERT), enabling the Transformer-based generator to more accurately model the nuanced clinical patterns inherent in real EHRs. Extensive experiments on the MIMIC-III and MIMIC-IV datasets demonstrate that HiSGT significantly improves the statistical alignment of synthetic data with real patient records, as well as supports robust downstream applications such as chronic disease classification. The code is available at https://github.com/jameszhou-gl/HiSGT. .
[…]
4194
@@ 1 Introduction
Electronic Health Records (EHRs) are a cornerstone of modern healthcare, playing a pivotal role in patient care, clinical decision-making, and medical research [17, 8, 38, 39]. Despite their importance, strict regulations surrounding privacy and data confidentiality have made it challenging to share and utilize EHR data at scale [18, 19]. As a result, there is growing interest in synthetic EHR data generation, a paradigm that aims to produce clinically realistic datasets mimicking the distributional properties of real EHRs while protecting sensitive patient information [14, 25, 35, 12].
Recent advances in generative models have spurred significant progress in synthetic EHR data generation. Many approaches use Generative Adversarial Networks (GANs) [7, 44, 34, 9, 43, 30, 21] and Variational AutoEncoders (VAEs) [4]. Other methods have drawn on the Transformer architecture [36] to capture long-range dependencies in clinical sequences. In these models, medical codes and clinical events are typically treated as discrete tokens, with a next-token prediction objective used to model the sequential structure of EHR data [33, 20, 28, 31]. However, they overlook the rich semantic information embedded in clinical code descriptions and the inherent hierarchical relationships in clinical coding systems. Consequently, the synthesized patient sequences frequently lack high clinical fidelity, which limits the utility of the generated EHR data. In particular, two major research gaps remain.
First, medical coding systems such as the International Classification of Diseases (ICD) [27] are organized hierarchically, with codes exhibiting parent–child and sibling relationships that reflect key clinical knowledge. For example, in ICD-10, the code ‘E11’ (Type 2 diabetes mellitus) serves as a parent for ‘E11.6’ (type 2 diabetes mellitus with other specified complications) and as an ancestor for ‘E11.65’ (Type 2 diabetes mellitus with hyperglycemia), while sibling codes such as ‘E11.65’ and ‘E11.64’ (Type 2 diabetes mellitus with hypoglycemia) capture different diabetes complications at a more granular level. In clinical practice, a patient may first be diagnosed at a coarse level with ‘E11’ and later be assessed for specific complications (e.g., ‘E11.64’ or ‘E11.65’). However, current generative models often treat each code as an isolated symbol, failing to leverage these multi-level taxonomic properties. As a result, they may generate conflicting sibling codes that cannot clinically co-occur (e.g., ‘E11.64’ hypoglycemia and ‘E11.65’ hyperglycemia in the same visit) or neglect important subcodes.
Second, each medical code is accompanied by a textual description that clarifies its clinical semantics and distinguishes between disease subtypes. For instance, ‘E11.65’ denotes ‘Type 2 diabetes mellitus with hyperglycemia’, while ‘Z13.6’ specifies ‘screening for cardiovascular disorders’. Such text descriptions are essential for determining the precise clinical focus of a patient’s visit. In practice, each patient’s visit to the hospital often centers on a primary diagnosis (or a small set of diagnoses) with additional procedures tailored to that theme until the patient’s condition is confirmed by the physician. Consequently, the codes assigned within a single visit tend to be semantically related, reflecting the shared focus on a particular disease process. However, existing generative methods typically disregard these semantic descriptions and treat codes solely as discrete tokens. Without leveraging the underlying semantics, these models may synthesize codes that are logically or clinically incompatible.
@@
To address these limitations, we propose a novel framework called Hierarchy and Semantics-Guided Transformer (HiSGT), which integrates both hierarchical relationships and semantic information into the generative process. As shown in Figure 1, HiGST begins by constructing a hierarchical graph to mirror the taxonomy of real coding systems such as ICD, preserving the parent–child and sibling relationships. This graph helps to encode and extract clinically coherent paths for every single code. A graph neural network (GNN) is then employed to learn hierarchy-aware embeddings that capture structured dependencies among medical codes. Additionally, HiGST enriches each medical code with semantic embeddings derived from a pre-trained clinical language model (e.g. ClinicalBERT [2]). These embeddings are consistent with their text description to help inject essential domain knowledge into the Transformer backbone. Thus, concepts like ‘Type 2 diabetes’ carry meaningful distinctions rather than being treated solely as a meaningless token. Finally, these refined embeddings are fed into a Transformer-based generative model to synthesize EHR data that are both semantically consistent and structurally coherent.
Our work makes the following key contributions to the field of synthetic EHR generation: (1) we propose a novel framework that leverages both hierarchical graph representations and clinical semantic embeddings, as an approach to embed medical codes in general, to overcome the limitations of raw code-only generative methods; (2) we develop an enhanced Transformer model that integrates refined code representations, enabling the generation of synthetic EHRs with high clinical fidelity; (3) we demonstrate the effectiveness of HiSGT on two real-world EHR datasets, showing significant improvements in synthetic data fidelity, utility for downstream tasks (e.g., disease classification), and privacy preservation.
2 Related Work
2.1 EHR Data Generation
The generation of synthetic EHRs have evolved across several categories [6]. Eearly rule-based methods rely on predefined clinical guidelines to simulate patient records [5, 26]. While these methods ensure clinically interpretability, they fail to capture complex correlations in real-world EHR data. GAN-based approaches model the distribution of real EHR data through adversarial training. A generator synthesizes patient records, while a discriminator distinguishes between real and synthetic samples [43, 40, 37, 24, 3, 42, 22, 41, 34, 21]. VAE-based methods learn a latent representation of EHR data and generate new samples by sampling from the learned distribution [4, 32]. However, both of GANs and VAEs struggle with rare events in high-dimensional EHR data and capturing fine-grained longitudinal dependencies in EHR data. Recently, Transformer-based generative models gained traction due to their ability to capture long-range dependencies in patient records. HALO [33] applies a self-attention mechanisms to model the sequential structure of medical codes, significantly expanding the vocabulary of synthetic data from fewer than 100 tokens to over thousands. However, existing methods rely solely on modeling raw code sequences. As result, the generated records lack high clinical fidelity. In this work, we propose a Transformer-based framework with two inductive biases of hierarchical and semantic information.
2.2 Generative Transformer
The Transformer architecture, introduced by  [36], has revolutionized sequence modeling by effectively capturing long-range dependencies through self-attention mechanisms. In natural language processing (NLP), generative Transformer models such as GPTs series [29] have demonstrated SOTA performance in text generation. These models predict tokens sequentially while conditioning on all previous tokens, making them highly effective for learning complex dependencies. Beyond NLP, Transformers have been successfully adapted to other domains including computer vision [10] and multimodal learning [16]. In healthcare, Transformer-based models for synthetic EHR generation typically treat patient trajectories as sequences of discrete medical codes [33, 28]. Although they enable scaling to synthetic data with thousands of unique tokens, they model medical codes purely as discrete symbols, thereby the generated EHR lack clinical fidelity. Our work extends Transformer-based EHR generation by explicitly incorporating structured knowledge into the generative process—leveraging hierarchical representations and semantic embeddings inherent in clinical coding systems.
3 Methodology
3.1 Problem Formulation
EHR data naturally exhibits a multi-level structure comprising patients, visits and clinical events. Let $\mathcal{X}$, $\mathcal{V}$ and $\mathcal{E}$ denote the sets of patients, visits and clinical events, respectively. For each patient $\bm{x}\in\mathcal{X}$, the medical history is represented as a sequence of visits $\bm{x}=\{\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{n_{x}}\}$, where $n_{x}$ is the number of visits. Each visit $\bm{v}_{i}$ comprises a set of medical codes representing diagnoses, procedures, and other clinical events: $\bm{v}_{i}=\{e^{i}_{1},e^{i}_{2},\ldots,e^{i}_{m_{i}}\},$ with $m_{i}$ denoting the number of medical codes recorded during visit $\bm{v}_{i}$. The objective of synthetic EHR generation is to model the joint distribution of patient trajectories while preserving both intra-visit (within-visit) and inter-visit (longitudinal) dependencies. The model estimates the conditional probability of each clinical event given past visits and previously observed events within the same visit:
where $\Theta$ represents the model parameters. The generative model seeks to maximize this likelihood to ensure that the generated sequences closely resemble real-world clinical patterns. Once trained, this model generates realistic EHR trajectories auto-regressively. At each step $j$, the predicted event $e_{j}^{i}$ is sampled from the conditional distribution $P\Bigl{(}e^{i}_{j}\,\Big{|}\,\bm{v}_{1},\cdots,\bm{v}_{i-1},e^{i}_{1},\cdots,e^{i}_{j-1}\Bigr{)}$, appended to the sequence, and used as input for the subsequent step. This process repeats until a predefined maximum length $L$ is reached or an ‘END_RECORD’ token is generated, signaling completion.
3.2 HiSGT Framework
3.2.1 Tokenization
The tokenization process in HiSGT structures raw EHR data into a sequence of discrete tokens while preserving temporal dependencies betweeen medical events. We extract longitudinal patient records from the MIMIC-III and MIMIC-IV datasets, using key tables such as patients.csv, admissions.csv and diagnoses_icd.csv. Each patient’s health timeline consists of one or more hospital visits, with each visit containing diagnosis codes that describe the patient’s conditions. To ensure an accurate representation of patient trajectories, the raw data undergoes preprocessing, where hospital visits are first sorted chronologically to maintain the correct sequence of admissions. Diagnosis codes are then grouped by patient ID, ensuring that all events for a single patient remain together. Additionally, each patient is assigned phenotype labels derived from their diagnoses, according to predefined clinical mappings. Once preprocessed, the dataset is tokenized by constructing a vocabulary that includes all unique ICD codes, along with phenotype label tokens and special tokens.
As shown in Figure 1(a), each patient sequence follows a structured format. It begins with a START_RECORD token, marking the initialization of a new patient record. This is followed by phenotype label tokens that summarize the patient’s conditions, such as Hypertension with complications and Pneumonia and Chronic kidney disease. These label tokens play a crucial role in conditioning the generation of subsequent visits, ensuring consistency with standard utility evaluations. The END_LABEL token separates phenotype labels from individual hospital visits. Each visit is represented as a collection of medical event tokens corresponding to ICD codes, e.g., ICD9-40301 (Hypertensive chronic kidney disease) and ICD9-486 (Pneumonia, organism unspecified). To preserve temporal dependencies, events from the same visit are grouped and concluded with an END_VISIT token. The sequence may progress through multiple visits, maintaining the correct ordering of hospital admissions. Finally, the END_RECORD token indicates the end of a patient’s medical history. Since patient trajectories vary in length, PADDING tokens are appended where necessary to ensure consistent input dimensions.
3.2.2 Hierarchical & Semantic Embeddings
Semantic Embeddings. Beyond hierarchical relationships, medical codes are accompanied by textual descriptions that provide additional clinical context. These descriptions are essential for distinguishing between related conditions, procedures, and medications. For example, the ICD-10 code ‘E11.65’ specifies ‘Type 2 diabetes mellitus with hyperglycemia’ whereas ‘Z13.6’ specifies ‘screening for cardiovasscular disorders’. Ignoring these textual descriptions can lead to synthetically generated records that fail to reflect meaningful clinical distinctions. HiSGT employs ClinicalBERT [2], a pre-trained clinical language model, to encode these descriptions into dense semantic representations. Given the text description $T(e)$ corresponding to medical code $e$, ClinicalBERT generates a semantic embedding: $\bm{s}_{e}=\mathrm{BERT}(T(e))$. Since ClinicalBERT outputs embeddings in a fix and high-dimensional space, we apply a linear transformation to project them into the model’s latent space: $\bm{Z}_{s}=\bm{W}_{s}\bm{s}_{e}+\bm{b}_{s}$; where $\bm{W}_{s}\in\mathbb{R}^{d_{s}\times d}$ and $\bm{b}_{s}\in\mathbb{R}^{d}$ are learnable parameters.
$\bm{Z}\;=\;\bm{Z}_{t}\;+\;\bm{Z}_{h}\;+\;\bm{Z}_{s}$ (2)
3.2.3 Model Structure
HiSGT employs a decoder-only Transformer architecture [29] for auto-regressive (AR) generation of EHR sequences. As shown in Figure 1(c), the input consists of standard token embedding $\bm{Z}_{t}$ with pre-computed hierarchical ($\bm{Z}_{h}$) and semantic ($\bm{Z}_{s}$) embeddings, plus positional embeddings. This composite embedding is processed through six stacked Transformer decoder blocks using causal masking for realistic sequence generation. The final hidden stages are projected to three output heads: 1) a code head for next-token prediction via probabilities over the medical code vocabulary, 2) a hierarchical consistency head, and 3) s semantic consistency head. The latter two heads act as regularization, enforcing alignment with the input hierarchical and semantic structures (derived from ICD taxonomy and ClinicalBERT descriptions, respectively) to ensure clinically meaningful and structured output.
3.2.4 Training and Inference
HiSGT training uses a multi-objective approach, balancing a standard cross-entropy loss for next-token prediction with two auxiliary Mean Squared Error (MSE) losses from the consistency heads. These auxiliary losses act as implicit regularization mechanisms, guiding the model toward generating sequences that maintain both structural coherence and semantic fidelity. For inference, HiSGT generates patient trajectories auto-regressively: starting with a START_RECORD token, it predicts phenotype labels until END_LABEL, then generates sequences of medical codes per visit until an END_RECORD token is produced, completing the synthetic patient record.
4 Experiments
@@
Table 2: Performance on Phenotype Classification. This table presents the results of the Train on Synthetic, Test on Real (TSTR) evaluation for phenotype classification. Models are trained on synthetic patient records and evaluated on real MIMIC-III and MIMIC-IV datasets. We report the average accuracy, precision, recall, and F1-score (with standard deviations) across 25 phenotype categories.
@@
4.1 Dataset and Baseline Details
We conduct experiments on two widely used real-world EHR datasets: MIMIC-III [18] and MIMIC-IV [19]. We follow the tokenization steps in §3.2.1. After preprocessing, MIMIC-III consists of 46,520 patients with 6,984 unique ICD diagnosis codes and 7,012 unique tokens, while MIMIC-IV includes 124,525 patients with 9,072 unique ICD codes and 9,102 unique tokens. Both datasets contain structured longitudinal patient records with multiple hospital visits, making them well-suited for evaluating synthetic EHR generation models. To assess the effectiveness of HiSGT, we compare it against six state-of-the-art (SOTA) generative models used for EHR synthesis. LSTM [23] is a recurrent neural network that models sequential dependencies. GPT [29] applies a transformer-based autoregressive approach for next-token prediction. EVA [4] uses a variational autoencoder (VAE) to learn latent representations of EHR sequences. SynTEG [44] employs a GAN-based framework to generate synthetic patient visits. HALO-Coarse [33] treats each visit as a single token, modeling longitudinal dependencies at a coarse level, while HALO [33] refines this approach by representing visits at a more granular level.
4.2 Experimental Setting
We conduct experiments using the MIMIC-III v1.4 and MIMIC-IV v2.2 datasets. The data is split into 80%-20% for training and testing, with an additional 90%-10% split within the training set for validation. For HiSGT, we use a six-layer Transformer architecture with eight attention heads and a hidden size of 384. Training is performed using the Adam optimizer with a learning rate of $10^{-4}$, a batch size of 48, and a dropout rate of 0.1. The input sequence length is set to 768 tokens, and training is run for 100 epochs with early stopping (patience = 10) based on validation loss. The best model checkpoint is saved according to minimum validation loss. The model is implemented in Python 3.13.0 using PyTorch 2.5.0, along with scikit-learn 1.5.2, NumPy 2.1.2, and transformers 4.46.2. All experiments are conducted on a single NVIDIA H100 GPU (80GB VRAM).
4.3 Main Results
4.3.1 Fidelity Evaluation
To assess the clinical fidelity of synthetic EHR data, we employ four complementary metrics that assess the alignment between real and synthetic records using the $R^{2}$ coefficient. These capture individual event distributions and relationships within and across visits. First, the Unigram score measures how well the marginal distribution of individual medical events is preserved by comparing event frequencies. Beyond individual events, the Bigram score evaluates intra-visit coherence by capturing statistical dependencies between pairs of medical events within the same visit—an essential property, as these events often shown strong correlations. Similarly, the (Seq Bigram) score extends this to inter-visit dependencies, quantifying how well longitudinal patterns (transitions between medical codes across consecutive visits) are captured. Finally, (DimWise) evaluates alignment at the patient level by comparing the mean per-patient probability distribution of medical codes. We compute the relative frequency of each code per patient, average these probabilities across all patients for each dataset (real and synthetic), and calculate the $R^{2}$ between the resulting vectors.
4.3.2 Utility Evaluation
Table 3: Phenotype Classification Performance Across Disease Categories. This table provides a detailed breakdown of phenotype classification performance for different disease types (acute, chronic, and mixed). We compare models trained on synthetic data (HALO and HiSGT) with real data as the ground truth. Accuracy and F1-scores are reported for each phenotype category, with macro-averaged scores included for all acute, mixed, and chronic disease groups.